OpenStudy (anonymous):

A man weighs 750 N on the surface of the earth. What would be his weight when standing on the moon? The masses of the earth and the moon are respectively, 5.98×1024 kg and 7.36×1022 kg. Their radii are respectively 6.37×103 km and 1.74×103 km

OpenStudy (anonymous):

Will use Newton’s Law of Gravity. F=(kG)(mM)/(D^2) k=.224 (is a constant that makes the force come out in pounds) G=6.67*10^-11 (Gravitational Constant) M=7.36*10^22kg (the mass of the Moon) m=76.478715kg (mass of the person) D=1.74x10^6 meters (the radius of the moon) F=(.224)(6.67*10^-11)(76.478715kg)(7.3*10^22)/(1.74*10^6)^2 = 27.55 lb that means he will weigh 27.55lb or 28 lb on the surface of the Moon. http://muller.lbl.gov/teaching/physics10/old%20physics%2010/chapters%20(old)/chapters_Jan_2004/Chapter03.pdf

OpenStudy (anonymous):

pls the options are :200.5N,123.7N,550.4N AND 1000.0N..i need an explanation

OpenStudy (anonymous):

expecting ur explanations,the options are in N not Ib..THANKS

OpenStudy (anonymous):

So the answer is 123.7N ( sorry if this doesn't help :( ) You could also convert that 27.55lb into N by using the converting factor 1lb=4.45N so you'll multiply 27.55*4.45N = 123N

OpenStudy (anonymous):

WAOO,tnx so much

OpenStudy (anonymous):

A ball of mass 50 g tied to the end of a 50 cm inextensible string is whirled around in a vertical circle. Find the tension in the string when the ball is at the top of the circle. Take g=10m/s2

OpenStudy (anonymous):

pls help me with that question